# Robin: a Suite of Multi-Scale Vision-Language Models and the CHIRP Evaluation Benchmark

### v2

@@ 6.4 Elo ratings
The results from the human and VLM evaluations using this average Elo rating is shown in Figure 5.
@@
Figure 6: Robin model performance on CHIRP’s “overall” criteria measured using different evaluators. Median Elo scores shown calculated over 1000 bootstraps on the criteria. Left. Elo calculated via different evaluators on the LLM Size and VE Size ablation matchups. Right. Elo calculated from GPT-4V (R) and human survey across the entire suite.
6.5 Agreement
@@ 6.6 Observations and Insights
6.7 Limitations
Although the CHIRP benchmark revealed scaling trends in our models that other benchmarks did not, it has several notable limitations. First, it heavily relies on the strong language proficiency of the evaluator, to evaluate a models’ perceptual capabilities.
@@ 7 Conclusions
8 Acknowledgments
We acknowledge the support from the Canada CIFAR AI Chair Program and from the Canada Excellence Research Chairs (CERC) Program. This project used compute resources provided by the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. This project further used compute resources provided by Mila, and Compute Canada.
@@ A.3 Final loss plots
Figure 8: Heatmap showing the loss of the different Robin models (lower is better).
Figure 9: Log-log plot showing how the loss scales with total parameter count.
@@ B.2 Verifying the results on an independent SoTA model
While other experiments in this paper are done using our Robin scaling suite with based on the Pythia LLMs, this section was done using LLaVA1.5-7B Liu et al. (2023a), in order to make sure that our results translated across models. We manually graded LLaVA1.5-7B responses on our sample of GQA questions. Using the known evaluations, we graded the accuracy of automated, LLM, and VLM evaluation methods. Results of grading accuracy on LLaVA1.5-7B are presented in the confusion matrix 13.
@@ Appendix C CHIRP
Benchmark questions and images available at https://huggingface.co/datasets/cerc-aai/CHIRP
@@ C.2.2 Human Survey
[…]
@@ C.2.3 AI evaluations
AI evaluations were run on all 10 matchups for both the LLM size study and VE size study. For the all Robin models study, we only ran GPT-4V (R) on a sample of 50 of the 190 possible matchups for each question. The prompts used for the evaluation are shown in Figures 20 and 21.
category_prompts = [ "Which response do you prefer overall, considering all factors?", "Which response is more relevant to the question and provides a more complete answer?", "Which response displays a better understanding of concepts and better reasoning?", "Which response more accurately describes the image without adding or describing objects or elements that dont exist in the image?", "Which responses description of the image is more detailed, taking into consideration both the amount and quality of the details provided?" ] response_eval = f""" Here are two responses to the same question: Response 1: {response_1} Response 2: {response_2} {category_prompts[category]} Respond with the number 1 or 2 corresponding to the better answer. """ messages = [ {"role": "user", "content": <IMAGE> + f"{question}"} {"role": "llava", "content": LLaVAs_response} {"role": "user", "content": response_eval} ]

### v3

@@ 6.1 Qualitative Analysis - Trends
The results from the human and VLM evaluations using average Elo rating is shown in Figure 5.
@@
6.2 Quantitative Analysis - Agreement
@@ 6.3 Observations and Insights
7 Limitations
Although the CHIRP benchmark revealed scaling trends in our models that other benchmarks did not, it has several notable limitations. First, it heavily relies on the strong language proficiency of the evaluator, to evaluate a models’ perceptual capabilities. This is particularly relevant to frontier AI models, which will require human evaluations.
@@ Appendix A Model training setup
Model Parameter Count
CLIP ViT B 86 million
CLIP ViT L 307 million
CLIP ViT H 632 million
CLIP ViT g 1 billion
@@ A.3 Final loss plots
(a) Heatmap showing the loss of the different Robin models.
(b) Log-log plot showing how the loss scales with total parameter count.
Figure 8: Comparison of loss scaling and evaluation scores for the Robin models.
@@ A.4 Detailed benchmark scores of all the Robin scaling suite models
A.5 Additional benchmarks
Table 6: List of the results obtained by every model of the Robin scaling suite.
LLM VE OCR Bench MathVista
Pythia 410M CLIP ViT B 16 1.1% 20.6%
Pythia 410M CLIP ViT L 14 0.7% 21.2%
Pythia 410M CLIP ViT H 14 0.8% 17.4%
Pythia 410M CLIP ViT g 14 0.1% 20.1%
Pythia 1.4B CLIP ViT B 16 10.4% 20.2%
Pythia 1.4B CLIP ViT L 14 10.3% 20.0%
Pythia 1.4B CLIP ViT H 14 8.2% 18.1%
Pythia 1.4B CLIP ViT g 14 6.5% 19.9%
Pythia 2.8B CLIP ViT B 16 8.7% 20.0%
Pythia 2.8B CLIP ViT L 14 0.7% 17.7%
Pythia 2.8B CLIP ViT H 14 4.1% 17.3%
Pythia 2.8B CLIP ViT g 14 6.1% 19.2%
Pythia 6.9B CLIP ViT B 16 10.2% 21.8%
Pythia 6.9B CLIP ViT L 14 16.7% 23.1%
Pythia 6.9B CLIP ViT H 14 10.5% 17.0%
Pythia 6.9B CLIP ViT g 14 9.2% 20.2%
Pythia 12B CLIP ViT B 16 1.6% 19.7%
Pythia 12B CLIP ViT L 14 10.6% 19.5%
Pythia 12B CLIP ViT H 14 9.1% 20.1%
Pythia 12B CLIP ViT g 14 7.9% 20.4%
@@ B.2 Verifying the results on an independent SoTA model
While other experiments in this paper are done using our Robin scaling suite with based on the Pythia LLMs, this section was done using LLaVA1.5-7B Liu et al. (2023a), in order to make sure that our results translated across models. We manually graded LLaVA1.5-7B responses on our sample of GQA questions. Using the known evaluations, we graded the accuracy of automated, LLM, and VLM evaluation methods. Results of grading accuracy on LLaVA1.5-7B are presented in the confusion matrix 14.
@@ Appendix C CHIRP
Benchmark questions and images available at https://huggingface.co/datasets/Anonymous1234565/CHIRP
@@ C.2.2 Human Survey
For the LLM size study, 5 of the 10 combinations of matchups between models were randomly sampled for each question. For each of those sampled matchups, we asked participants to indicate their preference for the 5 evaluation criteria. We only allowed each participant to respond with their preferences for a single matchup. This led to a total of 520 individual responses ($104\text{ questions}*5\text{ matchups}$). A breakdown of the questions asked in each survey is presented in Table 15. An example of the participant interface is shown in Figure 20.
@@ C.2.3 AI evaluations
AI evaluations were run on all 10 matchups for both the LLM size study and VE size study. For the all Robin models study, we only ran GPT-4V (R) on a sample of 50 of the 190 possible matchups for each question. The prompts used for the evaluation are shown in Figures 21 and 22.
category_prompts = [ "Whichresponsedoyoupreferoverall,consideringallfactors?", "Whichresponseismorerelevanttothequestionandprovidesamorecompleteanswer?", "Whichresponsedisplaysabetterunderstandingofconceptsandbetterreasoning?", "Whichresponsemoreaccuratelydescribestheimagewithoutaddingordescribingobjectsorelementsthatdon’texistintheimage?", "Whichresponse’sdescriptionoftheimageismoredetailed,takingintoconsiderationboththeamountandqualityofthedetailsprovided?" ] response_eval = f""" Herearetworesponsestothesamequestion: Response1:{response_1} Response2:{response_2} {category_prompts[category]} Respondwiththenumber1or2correspondingtothebetteranswer. """ messages = [ {"role": "user", "content": <IMAGE> + f"{question}"} {"role": "llava", "content": LLaVAs_response} {"role": "user", "content": response_eval} ]